OpenStudy (anonymous):

Which line is perpendicular to the line y = x + 11 ?

OpenStudy (anonymous):

y=-x+11 I believe...

OpenStudy (triciaal):

slope of the given line = 1 and when x = 0 y = 11 so it passes through (0,11) slope of perpendicular = -1/m = -1 and passes through (0, 11) another point on the perpendicular = (x, y) using point-slope formula for the equation of the perpendicular line , (y - 11) = -(x - 0) y = -x + 11 yes